# Optimize Weight Rounding via Signed Gradient Descent for the Quantization of LLMs

### v4

@@ Abstract
Large Language Models (LLMs) have demonstrated exceptional proficiency in language-related tasks, but their deployment poses significant challenges due to substantial memory and storage requirements. Weight-only quantization has emerged as a promising solution to address these challenges. Previous research suggests that fine-tuning through up and down rounding can enhance performance. In this study, we introduce SignRound, a method that utilizes signed gradient descent (SignSGD) to optimize rounding values and weight clipping within just 200 steps. SignRound integrates the advantages of Quantization-Aware Training (QAT) and Post-Training Quantization (PTQ), achieving exceptional results across 2 to 4 bits while maintaining low tuning costs and avoiding additional inference overhead. For example, SignRound achieves absolute average accuracy improvements ranging from 6.91% to 33.22% at 2 bits. It also demonstrates robust generalization to recent models and achieves near-lossless quantization in most scenarios at 4 bits. The source code is publicly available at https://github.com/intel/auto-round.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable proficiency in language-related tasks (OpenAI, 2022; Touvron et al., 2023a). However, deploying LLMs poses significant challenges due to their extensive memory and storage requirements. Additionally, the computational demands of these models create obstacles for real-time applications. Thus, studying techniques such as quantization becomes crucial to enable the efficient deployment of LLMs. Quantization techniques can be broadly categorized into two main types: quantization-aware training (QAT) (Esser et al., 2019; Zhuang et al., 2021; Lee et al., 2021; Liu et al., 2023b) and post-training quantization (PTQ) (Nagel et al., 2019; Xiao et al., 2022; Frantar et al., 2022; Nagel et al., 2020).
QAT involves training the model with quantization in mind. During QAT, the model is trained using simulated lower-precision representations, allowing it to learn and adapt to the effects of quantization. This approach often results in better accuracy compared to PTQ. However, QAT does come with certain drawbacks, including increased training complexity, longer training times, and the need to tune hyperparameters. The application of QAT to LLMs can be particularly costly, despite recent efforts (Hu et al., 2021; Dettmers et al., 2023) to improve the efficiency of fine-tuning LLMs.
@@
Two types of tensors commonly undergo quantization: activations and weights. Weight-only quantization has recently gained traction for LLMs due to its advantageous tradeoff. Quantizing activations for LLMs can be complex (Wei et al., 2022b; Xiao et al., 2023; Bondarenko et al., 2023), making weight-only quantization a more practical and feasible option. Moreover, the main bottleneck in generating new tokens for LLMs often arises from memory bandwidth limitations (Kim et al., 2023a), emphasizing the relevance of weight-only quantization.
@@
This study focuses on weight-only quantization. In quantizing weights, a critical step involves rounding, primarily achieved through rounding-to-nearest (RTN). RTN quantizes each weight independently by rounding it to the nearest integer. However, RTN overlooks the relationships between weights as well as between weights and activations. Nagel et al. (Nagel et al., 2020) explored the potential for an enhanced rounding strategy to improve accuracy. They approached the rounding task by formulating it as a quadratic unconstrained binary optimization problem and approximating the loss using a Taylor series expansion. However, relying solely on the second-order term may not yield accurate results, as rounding can significantly modify weights, making other order terms non-negligible.
We chose SignSGD as our optimization method to approach the optimal rounding solution within a limited number of steps, drawing inspiration from the well-defined boundaries of the solution space. This space is confined to ranges of [-0.5, 0.5] for rounding and [0, 1] for weight clip scales, offering several advantages for SignSGD. Firstly, the optimal value for up and down rounding typically resides in a large region rather than a single float, as only the threshold for altering the rounding value is significant. This eliminates the necessity for the gradient magnitude to converge precisely to a single point. Secondly, due to these confined boundaries, SignSGD allows efficient navigation of this space within a limited number of steps. In contrast, optimizers like Adam (Kingma and Ba, 2014) may struggle due to significant variations in gradient magnitude, making it challenging to converge to the optimal value within a restricted number of steps. Thirdly, SignSGD is inherently intuitive, facilitating easy adjustment of the step size (learning rate). For example, we employed the same optimizer hyperparameters across all experiments unless explicitly stated, consisting of 200 steps and a learning rate of $35e-$ 3, with linear weight decay. This ensures that $200*0.005/2=0.5$ covers the range of [-0.5, 0.5] for rounding and [0.5, 1] for weight clipping, which works well in practice. Fourthly, SignSGD stands out for its lightweight nature compared to other optimizers, requiring less memory and computational resources. Figure 1 provides an overview of our method. Our contributions primarily lie in three aspects:
@@
  • We introduce a concise yet effective method for optimizing the weight only quantization, combining the strengths of both QAT and PTQ. Our approach leverages SignSGD to tune the rounding with the weight clip, without introducing any additional overhead during inference.
@@
  • Our empirical results demonstrate a notable or significant enhancement in performance compared to recent works across different quantization configurations, from 2-bit to 4-bit.
@@
  • We demonstrate that SignRound’s performance can be further enhanced by fine-tuning model-wise hyperparameters within a constrained space. Moreover, our method demonstrates strong generalization across various models and delivers nearly lossless results across the majority of scenarios using 4-bit quantization.
@@ 2 Related Work
QAT methods have gained widespread popularity in model compression, as they enable the fine-tuning process  (Esser et al., 2019; Zhuang et al., 2021; Lee et al., 2021)  (Zhuang et al., 2021), often leading to superior accuracy compared to the PTQ method,
@@
PTQ methods simplify the quantization process without the needs of additional training. (Nagel et al., 2019; Liu et al., 2021; Frantar and Alistarh, 2022; Hassibi et al., 1993; Yao et al., 2021; Wang et al., 2019). Given its low resource requirement, PTQ is particularly suitable for the quantization of Large Language Models (LLMs).
@@
Significant strides have been made in addressing the pressing need for quantizing large language models (LLMs). LLM.int8() (Dettmers et al., 2022) introduces a mixed-precision approach to preserve crucial channels in high precision. PEQA (Kim et al., 2024) presents a simple yet effective method that combines the benefits of parameter-efficient fine-tuning (PEFT) with quantized LLMs. AQLM (Mao et al., 2024) builds upon Additive Quantization, a classic algorithm from the Multi-Codebook Quantization family, adapting it to LLM quantization. ZeroQuantV2 (Yao et al., 2023) employs low-rank matrices to enhance model quality recovery. RPTQ (Yuan et al., 2023) addresses range differences between channels by rearranging and quantizing them in clusters. Other methods, such as SPIQ (Yvinec et al., 2023b), SmoothQuant (Xiao et al., 2022), and Outlier Suppression+ (Wei et al., 2023), utilize handcrafted equivalent transformations to mitigate quantization errors. LLM-QAT (Liu et al., 2023b) employs QAT to enhance performance.
@@
Weight-only quantization reduces the memory footprint and bandwidth demands by quantizing only the weights while retaining activations in floating-point precision, offering a promising balance between accuracy and compression. GPTQ (Frantar et al., 2022) optimizes weights using the Optimal Brain Surgeon technique (Hassibi et al., 1993), achieving low-bit quantization on LLMs with minimal tuning overhead. AWQ (Lin et al., 2023) follows the equivalent transformation approach with additional tuning in a constrained space, sharing similar limitations with SmoothQuant (Xiao et al., 2022). TEQ(Cheng et al., 2023) and OmniQuant (Shao et al., 2023) both utilize a trainable equivalent transformation, while OmniQuant employs extra weight clip tuning. HQQ (Badri and Shaji, 2023) accelerates quantization for large models by eliminating the need for calibration data, making the quantization process extremely fast. Some other works have incorporated optimization methods with extra inference overhead to improve quantization accuracy, such as dense-and-sparse decomposition techniques in SqueezeLLM (Kim et al., 2023a) and EasyQuant (Tang et al., 2024), as well as nonuniform quantization methods in NUPES (Yvinec et al., 2023a), QuIP#(Tseng et al., 2024),(Gong et al., 2024), AQLM(Mao et al., 2024), etc. Additionally, FineQuant (Kim et al., 2023b) introduces a straightforward heuristic weight quantization approach that adaptively determines quantization granularity. In this work, we focus on approaches that do not introduce overhead during inference.
@@
Adaptive Rounding (Nagel et al., 2020) has already showcased the potential of an advanced rounding strategy to enhance accuracy (Li et al., 2021; Wei et al., 2022a). They used the rounding task as a quadratic unconstrained binary optimization problem by approximating the task loss through a Taylor series expansion. However, considering only the second-order term may not yield accurate results. This is because the rounding value gets multiplied by a scaling coefficient during de-quantization, potentially introducing significant weight changes that make other order terms non-negligible. FlexRound (Lee et al., 2023) introduces a more flexible approach to rounding by incorporating element-wise division. However, it’s not easily scalable to apply to LLMs due to the needs of specialized hyperparameters for each specific model and task. Furthermore, Oscillation-free (Liu et al., 2023a) suggests that the introduction of learnable parameters might result in weight oscillation problems. AQuant (Li et al., 2022) introduced a dynamic approach where the border becomes a function dependent on the activation value to reduce the quantization error of activation.
@@
Signed gradient descent is not commonly utilized and is typically applied in specific scenarios, such as reducing communication costs. This is because signed gradient carries significantly less information compared to original gradient. Recent studies have shed light on the advantages of sign-based methods over gradient descent in certain conditions. Safaryan et al. (Safaryan and Richtárik, 2021) found that sign-based methods are preferable when the Hessian matrix is concentrated on its diagonal and the maximal eigenvalue is much larger than the average eigenvalue. Li et al. (Li et al., 2023a) investigated a variant of sign-based gradient descent that exhibits faster convergence. Additionally, Safaryan et al. (Safaryan and Richtárik, 2021) proposed a stochastic sign descent with momentum, which converges under the standard bounded variance assumption with the optimal asymptotic rate. These findings contribute to a better understanding of the potential benefits and applications of signed gradient descent methods.
Figure 1: An illustration of SignRound. Unlike the direct rounding in RTN, SignRound performs signed gradient descent to fine-tune the up and down rounding through block-wise output reconstruction. After lightweight forward and backward steps, WINT4 has been well optimized towards the minimal loss, therefore ready for the final inference deployment. Note that Quant and Dequant are two standard operations for quantization and dequantization respectively.
@@ 3 Methodology
We begin with an overview of quantization before delving into the specifics of our approach. The following operations can be utilized to quantize and de-quantize the weights $W$:
$\widetilde{W}=s*clip(\left\lfloor\frac{W}{s}+zp\right\rceil,n,m),n,m\in\mathbb{N}$ (1)
@@
$s=\frac{max(W)-min(W)}{2^{bit}-1}$ (2)
In order to improve the efficacy of the rounding quantization operation, we build upon prior research (Nagel et al., 2020) by introducing a single trainable parameter $V$ to adjust the rounding values.
$\widetilde{W}=s*clip(\left\lfloor\frac{W}{s}+zp+V\right\rceil,n,m),n,m\in\mathbb{N}$ (3)
$s=\frac{max(W)*\alpha-min(W)*\beta}{2^{bit}-1}$ (4)
$\min_{\alpha,\beta,V}\|WX-\widetilde{W}X\|_{F}^{2}$ (5)
Our method distinguishes itself primarily by leveraging SignSGD, with the motivation thoroughly outlined in Section Introduction 1. Figure 1 provides an illustration of our approach. And the pseudocode 1 presents more details of SignRound.
@@
3:     $d\leftarrow\textit{draw $bs$ samples from privious quantized block}$
@@
4 Experiments
This section presents a comprehensive evaluation of SignRound from various perspectives. We begin with a concise overview of the LLM architectures and tasks included in our assessment. Next, we provide a detailed comparison between our method and several existing approaches, emphasizing the unique advantages of SignRound. Furthermore, we conduct ablation studies to reinforce the efficacy of our choices and investigate the sensitivity of hyperparameters. Lastly, we evaluate the generation ability of our method across various recent models. The tuning cost comparisons are provided in Appendix A.
4.1 Experimental Settings
Evaluation and Tasks.
We evaluate multiple language tasks to address the task-agnostic setting. Specifically, we present the average accuracy results for 11 zero-shot tasks, including HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), PIQA (Bisk et al., 2020), LAMBADA (Paperno et al., 2016), TruthfulQA (Lin et al., 2021), OpenBookQA (Mihaylov et al., 2018), BoolQ (Clark et al., 2019), RTE (Dagan et al., 2010), ARC-Easy, ARC-Challenge (Clark et al., 2018), and MMLU (Hendrycks et al., 2020). We use lm-eval-harness (Gao et al., 2023) for all the above tasks. Furthermore, we complement our evaluation with perplexity (PPL) analysis on Wikitext2 (Merity et al., 2016), PTB (Marcus et al., 1993), and C4 (Raffel et al., 2020), following the source code [^1] https://github.com/IST-DASLab/gptq of GPTQ and Wikitext2 (Merity et al., 2016) using lm-eval-harness (Gao et al., 2023). However, we argue that perplexity is notably influenced by outliers, as illustrated in Table 14 for different algorithms. This susceptibility likely arises from the mathematical expression $\mathrm{PPL}(X)=\exp\left(-\frac{1}{t}\sum_{i=1}^{t}\log p_{\theta}(x_{i}|x_{<i})\right)$, where assigning a low probability to even one token can significantly inflate the perplexity score. Consequently, we prioritize the accuracy of the 11 tasks mentioned above as the primary metric, with perplexity data serving as supplementary reference.
Quantization Configurations.
In alignment with GPTQ (Frantar et al., 2022), our focus is specifically on weight-only quantization, targeting the linear layers within transformer blocks. Layers such as the embedding layer and typically the last linear layer like ’lm-head’ are excluded from the quantization process. Our evaluation primarily centers on W4, W4G128, and W3G128 configurations, where W4 indicates quantizing weights with 4 bits and G represents finer-grained grouping as described in (Park et al., 2022; Frantar et al., 2022). We adopt asymmetric quantization. To mitigate overfitting on the WikiText and C4 datasets, we randomly select 512 samples with the same seed from the readily available pile-10k dataset [^2] https://huggingface.co/datasets/NeelNanda/pile-10k for calibration, which comprises the first 10k samples from pile (Gao et al., 2020). We used a sequence length of 2048 for calibration, while for other methods, we adhere to their official settings.
Large Language Models.
Table 1: Average accuracies ($\uparrow$) across 11 tasks, as detailed in Section 4.1, for LLaMA and Mistral models at W2-W4. Our* denotes the highest accuracy achieved among the 8 hyperparameter choices, outlined in Section 4.2, whereas for the 70B model, we tested only a few options.
@@ 4.1 Experimental Settings
Table 2: Detailed accuracies($\uparrow$) across 11 tasks(0-shot) of LLaMA and Mistral models at W4G-1. Our* denotes the highest accuracy achieved among the 8 hyperparameter choices, outlined in Section 4.2, whereas for the 70B model, we tested only a few options. Appendix C provides more detailed data.
Mistral-7B FP16 61.35 75.68 61.27 74.03 80.79 28.03 32.80 83.67 67.51 80.81 50.34 63.30

### v5

@@ Abstract
Large Language Models (LLMs) have demonstrated exceptional proficiency in language-related tasks, but their deployment poses significant challenges due to substantial memory and storage requirements. Weight-only quantization has emerged as a promising solution, significantly reducing memory and storage needs without sacrificing too much performance. In this study, we introduce SignRound, a method that leverages signed gradient descent (SignSGD) to optimize rounding values and weight clipping in just 200 steps. SignRound integrates the advantages of Quantization-Aware Training (QAT) and Post-Training Quantization (PTQ), delivering exceptional results across 2 to 4 bits while minimizing tuning costs and avoiding additional inference overhead. For example, SignRound achieved absolute average accuracy improvements ranging from 6.91% to 33.22% at 2 bits, as measured by the average zero-shot accuracy across 11 tasks. It also demonstrates strong generalization in recent models, achieving near-lossless 4-bit quantization in most scenarios. The source code is publicly available at https://github.com/intel/auto-round.
Optimize Weight Rounding via Signed Gradient Descent for the Quantization of LLMs
Wenhua Cheng thanks: Correspondence:wenhua.cheng@intel.com  and Weiwei Zhang  and Haihao Shen  and Yiyang Cai Xin He and Kaokao Lv and Yi Liu Intel
@@ 1 Introduction
In recent years, there has been a significant surge in the adoption of Large Language Models (LLMs), leading to their widespread deployment demand even on devices with constrained resources. However, deploying LLMs on these devices poses significant challenges due to their extensive memory and storage requirements. Additionally, the computational demands of these models create obstacles for real-time applications. Therefore, studying techniques such as quantization is crucial for enabling the efficient deployment of LLMs. Quantization techniques can be broadly categorized into two main types: quantization-aware training (QAT) (Esser et al., 2020; Zhuang et al., 2021; Lee et al., 2021; Liu et al., 2023b) and post-training quantization (PTQ) (Nagel et al., 2019; Xiao et al., 2023; Frantar et al., 2022; Nagel et al., 2020).
QAT involves training the model with quantization in mind, using simulated lower-precision representations to allow the model to learn and adapt to the effects of quantization. This approach often results in better accuracy compared to PTQ. However, QAT has drawbacks, including increased training complexity, longer training times, and the need to tune hyperparameters. The application of QAT to LLMs can be particularly resource-intensive, despite recent efforts (Hu et al., 2021; Dettmers et al., 2023) to improve the efficiency of fine-tuning LLMs.
@@
Quantization commonly applies to two types of tensors: activations and weights. Quantizing activations for LLMs can be challenging (Wei et al., 2023; Xiao et al., 2023; Bondarenko et al., 2024), making weight-only quantization a more practical option. Moreover, the main bottleneck in generating new tokens for LLMs often arises from memory bandwidth limitations (Kim et al., 2023a), emphasizing the advantage of weight-only quantization.
@@
This study focuses on weight-only quantization. In quantizing weights, a critical step involves rounding, primarily achieved through rounding-to-nearest (RTN). RTN quantizes each weight independently by rounding it to the nearest integer, but it overlooks the relationships between weights and between weights and activations. Adaptive Rounding (Nagel et al., 2020) explored the potential for an enhanced rounding strategy to improve accuracy. They approached the rounding task by formulating it as a quadratic unconstrained binary optimization problem and approximating the loss using a Taylor series expansion. However, relying solely on the second-order term may not yield accurate results, as rounding can significantly modify weights, making other order terms non-negligible.
We select SignSGD(Balles et al., 2020; Li et al., 2023a; Safaryan and Richtárik, 2021) as our optimization method to approach the optimal rounding solution within a limited number of steps. The motivation behind this choice, which is elaborated in Section 3, stems from the well-defined boundaries of the solution space and the inherent simplicity of the method that necessitates only minimal hyperparameter tuning. Figure 1 provides an overview of our method. Our contributions primarily lie in three aspects:
@@
  • We introduce a concise yet effective method for optimizing the weight only quantization, combining the strengths of both QAT and PTQ. Our approach leverages SignSGD to tune the rounding with the weight clipping, without introducing any additional overhead during inference.
@@
  • Our empirical results demonstrate a significant performance enhancement compared to recent works across various quantization configurations, ranging from 2-bit to 4-bit.
@@
  • We demonstrate that SignRound’s performance can be further enhanced by fine-tuning model-specific hyperparameters within a constrained space. Moreover, our method demonstrates strong generalization across various models and delivers nearly lossless results across the majority of scenarios using 4-bit quantization.
Figure 1: An illustration of SignRound. Unlike the direct rounding in RTN, SignRound performs signed gradient descent to fine-tune the rounding and weight clipping through block-wise output reconstruction. After lightweight forward and backward steps, $\textbf{W}_{\text{INT4}}$ has been well optimized. Note that Quant and Dequant are two standard operations for quantization and dequantization respectively.
@@ 2 Related Work
QAT methods have gained widespread popularity in model compression, as they enable the fine-tuning process  (Esser et al., 2020; Zhuang et al., 2021; Lee et al., 2021), often leading to superior accuracy compared to the PTQ method.
@@
PTQ methods simplify the quantization process without the need for additional training. (Nagel et al., 2019; Liu et al., 2021; Frantar and Alistarh, 2022; Hassibi et al., 1993; Yao et al., 2021). Given its low resource requirement, PTQ is particularly suitable for the quantization of Large Language Models.
@@
Significant strides have been made in addressing the pressing need for quantizing large language models (LLMs). GPT3.int8() (Dettmers et al., 2022) introduces a mixed-precision approach to preserve crucial channels in high precision. AQLM (Mao et al., 2024) builds upon Additive Quantization, a classic algorithm from the Multi-Codebook Quantization family, adapting it to LLM quantization. ZeroQuantV2 (Yao et al., 2024) employs low-rank matrices to enhance model quality recovery. RPTQ (Yuan et al., 2023) addresses range differences between channels by rearranging and quantizing them in clusters. LLM-QAT (Liu et al., 2023b) employs QAT to enhance performance. Some other methods, such as SPIQ (Yvinec et al., 2023b), SmoothQuant (Xiao et al., 2023), and Outlier Suppression+ (Wei et al., 2023), utilize handcrafted equivalent transformations to mitigate quantization errors. These methods rely on the model architecture to fuse the equivalent transformation operations. LRQ (Lee et al., 2024) only needs to learn significantly fewer parameters while enabling the individual scaling of weights, thus boosting the generalization capability of quantized LLMs.
@@
Weight-only quantization reduces the memory footprint and bandwidth demands by quantizing only the weights while retaining activations in floating-point precision, offering a promising balance between accuracy and compression. GPTQ (Frantar et al., 2022) optimizes weights using the Optimal Brain Surgeon technique (Hassibi et al., 1993), achieving low-bit quantization on LLMs with minimal tuning overhead. AWQ (Lin et al., 2023) follows the equivalent transformation approach with additional tuning in a constrained space, sharing similar limitations with SmoothQuant (Xiao et al., 2023). TEQ (Cheng et al., 2023) and OmniQuant (Shao et al., 2023) both utilize a trainable equivalent transformation, while OmniQuant employs extra weight clip tuning. HQQ (Badri and Shaji, 2023) accelerates quantization for large models by eliminating the need for calibration data, making the quantization process extremely fast. Some other works have incorporated optimization methods with extra inference overhead to improve quantization accuracy, such as dense-and-sparse decomposition techniques in SqueezeLLM (Kim et al., 2023a) and EasyQuant (Tang et al., 2023), as well as nonuniform quantization methods in NUPES (Yvinec et al., 2023a), QuIP# (Tseng et al., 2024),(Gong et al., 2024), AQLM (Mao et al., 2024), etc. Additionally, FineQuant (Kim et al., 2023b) introduces a straightforward heuristic weight quantization approach that adaptively determines quantization granularity. In this work, we focus on approaches that do not introduce overhead during inference.
@@
Adaptive Rounding (Nagel et al., 2020) has already showcased the potential of an advanced rounding strategy to enhance accuracy (Li et al., 2021; Wei et al., 2022). They used the rounding task as a quadratic unconstrained binary optimization problem by approximating the task loss through a Taylor series expansion. However, considering only the second-order term may not yield accurate results. This is because the rounding value gets multiplied by a scaling coefficient during de-quantization, potentially introducing significant weight changes that make other order terms non-negligible. FlexRound (Lee et al., 2023) introduces a more flexible approach to rounding by incorporating element-wise division. However, it’s not easily scalable to apply to LLMs due to the needs of specialized hyperparameters for each specific model and task. Furthermore, Oscillation-free (Liu et al., 2023a) suggests that the introduction of learnable parameters might result in weight oscillation problems. AQuant (Li et al., 2022) introduced a dynamic approach where the border becomes a function dependent on the activation value to reduce the quantization error of activation.
@@
Signed gradient descent is not commonly utilized and is typically applied in specific scenarios, such as reducing communication costs. This is because signed gradient carries significantly less information compared to original gradient. Recent studies have shed light on the advantages of sign-based methods over gradient descent in certain conditions. Balles et al. (Balles et al., 2020) found that sign-based methods are preferable when the Hessian matrix is concentrated on its diagonal and the maximal eigenvalue is much larger than the average eigenvalue. Li et al. (Li et al., 2023a) investigated a variant of sign-based gradient descent that exhibits faster convergence. Safaryan et al. (Safaryan and Richtárik, 2021) proposed a stochastic sign descent with momentum, which converges under the standard bounded variance assumption with the optimal asymptotic rate. These findings contribute to a better understanding of the potential benefits and applications of signed gradient descent methods.
@@ 3 Methodology
We begin with an overview of quantization before delving into the specifics of our approach. The following operations can be utilized to quantize and dequantize the weights W:
$\mathbf{\widetilde{W}}=s*clip(\left\lfloor\frac{\mathbf{W}}{s}+zp\right\rceil,n,m),n,m\in\mathbb{N}$ (1)
@@
$s=\frac{max(\mathbf{W})-min(\mathbf{W})}{2^{bit}-1}$ (2)
@@
3:     $d\leftarrow\textit{draw $bs$ samples}$
@@
In order to improve the efficacy of the rounding quantization operation, we build upon prior research (Nagel et al., 2020) by introducing a trainable parameter $V$ to adjust the rounding values.
$\widetilde{\mathbf{W}}=s*clip(\left\lfloor\frac{\mathbf{W}}{s}+zp+\mathbf{V}\right\rceil,n,m),n,m\in\mathbb{N}$ (3)
$s=\frac{max(\mathbf{W})*\alpha-min(\mathbf{W})*\beta}{2^{bit}-1}$ (4)
$\min_{\alpha,\beta,\mathbf{V}}\|\mathbf{WX}-\mathbf{\widetilde{W}X}\|_{F}^{2}$ (5)
Our method distinguishes itself primarily by leveraging SignSGD, which optimizes parameters based on the sign of the gradients as follows:
$\mathbf{W}_{t+1}=\mathbf{W}_{t}-lr_{t}*sign(\mathbf{g}_{t})$ (6)
where t represents the step, lr is the learning rate and g denotes the gradient. The motivation is detailed below. Firstly, the optimal values for up and down rounding typically reside in a large region rather than a single float, as only the threshold for altering the rounding value is significant. This eliminates the necessity for the gradient magnitude to converge precisely to a single point. Secondly, due to the confined boundaries, i.e.[0.5,0.5]0.50.5[-0.5,0.5][ - 0.5 , 0.5 ] for rounding and $[0,1]$ for weight clipping, SignSGD allows efficient navigation of this space within a limited number of steps. In contrast, optimizers like Adam (Kingma and Ba, 2014) may struggle due to significant variations in gradient magnitude, making it challenging to converge to the optimal value within a restricted number of steps. Thirdly, SignSGD is inherently intuitive, facilitating easy adjustment of the step size (learning rate). For example, we employed the same optimizer hyperparameters across all experiments unless explicitly stated, consisting of 200 steps and a learning rate of 5e-3 with linear weight decay. Based on Eq. 6, the maximum adjustment for each parameter is the sum of the learning rates over all steps, that is, $200\times 0.005/2=0.5$. As a result,the adjustment can cover a range of [-0.5,0.5] when initialized at 0 for rounding, and a range of [0.5,1.0] when initialized at 1 and clipped to $\leq 1.0$ for weight clipping, which works well in practice. Fourth, SignSGD distinguishes itself by its lightweight design, demanding fewer memory and computational resources than optimizers like Adam(Kingma and Ba, 2014).
Figure 1 provides an illustration of our approach. And the Pseudocode 1 presents more details of SignRound.
@@
Table 1: Average accuracies ($\uparrow$) across 11 tasks, as detailed in Section 4.1, for LLaMA and Mistral models at W2-W4. ’Our*’ denotes the highest accuracy achieved among the 8 hyperparameter choices, outlined in Section 4.2, whereas for the 70B model, we tested only a few options.
4 Experiments
This section presents a comprehensive evaluation of SignRound from multiple perspectives. We begin with a brief overview of the LLM architectures and tasks included in our assessment. Next, we provide a detailed comparison between our method and several existing approaches, emphasizing the unique advantages of SignRound. Furthermore, we conduct ablation studies to reinforce the efficacy of our choices and investigate the sensitivity of hyperparameters. Lastly, we evaluate the generation ability of our method across various recent models. The tuning cost comparisons are provided in Appendix A.
4.1 Experimental Settings
Config Method Mmlu Lamb. Hella. Wino. Piqa Truth. Open. Boolq RTE ARC-e ARC-c. Avg.
W4G-1 16 bits 61.35 75.68 61.27 74.03 80.79 28.03 32.80 83.67 67.51 80.81 50.34 63.30